A 12 L gas kept in an expandable jar has an initial pressure of 2.1 atm at 25°C. If the jar is allowed to expand isothermally to 15 L, what will be the resulting pressure of the gas?
1
Expert's answer
2020-11-26T05:09:17-0500

pV=nRT.

n=pV/RT= 2.1 x 12 / 0.082 x 298 = 1,03126535 mol.

p=nRT/V=1,03126535 x 0.082 x 298/15 = 1,68000001 atm.